OpenStudy (anonymous):

sin[2tan(^-1)(5/12))

OpenStudy (anonymous):

Let \(\theta=\tan^{-1}\dfrac{5}{12}\). Then use the double angle identity: \[\sin2\theta=2\sin\theta\cos\theta=2\sin\left(\tan^{-1}\dfrac{5}{12}\right)\cos\left(\tan^{-1}\dfrac{5}{12}\right)\] From the substitution above, you have \(\tan\theta=\dfrac{5}{12}\), which comes from the following reference triangle:  Find the missing side with the Pythagorean theorem: \[5^2+12^2=x^2\] Knowing the value of \(x\) will allow you to determine the cosine and since of \(\theta\), since \[\cos\theta=\frac{12}{x}~~\text{and}~~\sin\theta=\frac{5}{x}\] So, \[\sin\left(2\tan^{-1}\dfrac{5}{12}\right)=2\left(\frac{5}{x}\right)\left(\frac{12}{x}\right)=\frac{120}{x^2}\]